OpenStudy (anonymous):

convert 1 dyne into newton

OpenStudy (michele_laino):

we have: \[1Newton = 1 Kg*1\frac{ m }{ s ^{2} }=10^{3}grams*10^{2}\frac{ cm }{ \sec ^{2} }=\] \[=10^{5}grams*\frac{ cm }{ \sec ^{2} }=10^{5}dyn\] so: \[1Newton=10^{5}dyn\] then \[1dyn=10^{-5}Newton\]